The electric field everywhere on the surface of a thin, spherical shell of radius 0.710 m is of magnitude 932 N/C and points radially toward the center of the sphere.     (a) What is the net charge within the sphere's surface?

ive calculated this out and i get 5.22E-8, but the computer system my school uses is saying thats wrong
Answer #1

let the charge be q.so,

E=kq/r^2

or 932=9*10^9*q/(0.71^2)

or q=5.22*10^-8 C

try changing the direction. use radially inwards(i guess the charge is negative)